OpenStudy (abbycross167):

can someone please help me with this math question? Please? I'll fan and give a medal!! Question - Answer the following question, indicating in words each step that is used. - A football field’s dimensions are represented by a width of (3x + 10) feet and a length of (7x + 10) feet. Find an expression for the area A of the football field. Give your answer as a quadratic trinomial.

OpenStudy (anonymous):

Double distribute. (3x+10)(7x+10) Multiply 3x by both terms in the second polynomial. So you get 21x^2 and 30x. Then take the 10 and multiply the it by the 7x and 10. Giving you 70x and 100. Combine like terms, you get 21x^2 +100x +100.Hope this helps.

OpenStudy (abbycross167):

This helps a lot!! thank you! I have to copy this down on Word, would you mind checking to make sure I copied them down right?

OpenStudy (abbycross167):

@mccarthym372390 here is how I wrote it down - (3x+10)(7x+10) 3x*7x=21x 3x*10=30x 10*7x=70x 10*10=100 21x^2+100x+100

OpenStudy (anonymous):

21x^2

OpenStudy (anonymous):

Also suggest writing it 3x*7x+3x*10 +7x*10+10*10. Unless you want to break it down.

OpenStudy (abbycross167):

oh ok so (3x+10)(7x+10) 3x*7x+3x*10 +10*7x+10*10 21x^2+30x+70x +100

OpenStudy (anonymous):

Yup!

OpenStudy (abbycross167):

great thank you so much!! :)